The E-JEM Company produces two types of laptop computer bags. The regular version requires $32 in capital, 4 hours of labor, and sells for S47. The deluxe version requires $38 in capital, 6 hours of labor, and sells for $57. How many of each type of bag should the company produce in order to maximize their revenue if they have $2000 and 280 labor hours available? The company should make regular laptop bags and » deluxe bags to maximize revenue.
Answer #1

To calculate regular bags and delux bags , we need to use simplex linear solver method . It is an optimization problem.

Where revenue maximizing equation = regular laptop bags* selling price of regular bags + delux laptop bags * selling price of delux bags

= 47*r + 57*d

Constraints will be as follows:

labor required for each regular bag * regular bags + labor required for each delux bag* delux bags<= labor limit

4*r + 6*d <= 280

Cost required for each regular bag * regular bags + cost required for each delux bag* delux bags<= cost limit

32*r + 38*d <= 2000

And r and d are non negative integers .

Using above equations in excel solver as follows:

after solving we get 34 regular bags and 24 delux bags with $2966 as maximum revenue .